I would want to know step by step of how to simplfy ecuation and to reduce the expresion?
1 4(2x-1)-6(x+3)
2 6a-2a+5=17
3 4(8-3w)=32-8(w+2)

Found 4 solutions by oscargut, checkley77, stanbon, Earlsdon:
Answer by oscargut(891)

1 4(2x-1)-6(x+3)=8x-4-6x-18=2x-22
2 6a-2a+5=17 then 4a+5=17 then 4a = 12 then a =3
3 4(8-3w)=32-8(w+2)then 32-12w=32-8w-16 then -12w+8w=32-16-32
then -4w=-16 then w=4

Answer by checkley77(12569)

1) 4(2x-1)-6(x+3)
8x-4-6x-18
2x-22
2(x-11)
-----------------------
2) 6a-2a+5=17
4a=17-5
4a=12
a=12/4
a=3
-----------------------
3) 4(8-3w)=32-8(w+2)
32-12w=32-8w-16
-12w+8w=-16
-4w=-16
w=-16/-4
w=4

Answer by stanbon(57282)

1 4(2x-1)-6(x+3)
8x-4 -6x-18
2x-22
------------------
2 6a-2a+5=17
4a = 12
a = 3
---------------
3 4(8-3w)=32-8(w+2)
32-12w = 32 -8w -16
-12w = -8w - 16
Add 12w to both sides; Add 16 to both sides to get:
4w = 16
w = 4
============
Cheers,
Stan H.

Answer by Earlsdon(6287)

Ok, the first one is not an ecuation(sic) but an expression.
1) 4(2x-1)-6(x+3) First, apply the distributive property to both terms to get:
(8x-4)-(6x+18) Now remove the parentheses and don't foget to change the sign (+ to -) in the second group.
8x-4-6x-18 Combine like terms.
(8x-6x)+(-4-18)
2x+(-22) = 2x-22 or 2(x-11)
------------------------------------------
2) 6a-2a+5 = 17 Now this is an equation! Combine the a's on the left side.
4a+5 = 17 Now subtract 5 from both sides.
4a = 12 Finally, divide both sides by 4.
a = 3
--------------------------------------------
3) 4(8-3w) = 32-8(w+2) Apply the distributive property, where appropriate.
(32-12w) = 32-(8w+16) Remove the parentheses and remember the sign-change (+ to -) in the right-most set of parentheses.
32-12w = 32-8w-16 Now combine like-terms.
32-12w = 16-8w Add 12 w to both sides.
32 = 16+4w Subtract 16 from both sides.
16 = 4w Finally, divide both sides by 4.
4 = w or w = 4